Calculate the future value of the following single amounts. (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use appropriate factor(s) from the tables provided. Round your answers to 2 decimal places.) Future Value Initial Investment 1. $ 8,000 6,000 | 3. 9,000 Annual Rate 10 % 12 % 8% Interest Period Compounded Invested Annu
Answer #1

calculate future value

Initial investment Annual rate Interest compounded Period invested Future value
1. 8000 10% Annually 7 years (8000*1.94872) = 15589.76
2. 6000 12% Semiannually 4 year (6000*1.59385) = 9563.10
5 9000 8% Quarterly 3 years (9000*1.26824) = 11414.16